OpenStudy (anonymous):

help please!!(12.5-3.2(6.4-3.5 times 1.4) +4.3) divided 6.

OpenStudy (anonymous):

\[\frac{ 12-3.2(6.4-3.5\times1.4)+4.3 }{ 6 }\] first solve bracket what is 3.5*1.4=

OpenStudy (anonymous):

3.5*1.4=4.9

OpenStudy (anonymous):

6.4-4.9=1.5

OpenStudy (anonymous):

3.2*1.5=4.8

OpenStudy (anonymous):

\[\frac{ 12-4.8+4.3 }{ 6 }=?\]

OpenStudy (anonymous):

0.48

OpenStudy (anonymous):

How you got 0.48?

OpenStudy (anonymous):

i do addition first and then i subtract by 12 and then i divide by 6.

OpenStudy (anonymous):

is it 1.916666667??

OpenStudy (anonymous):

\[=\frac{ 16.3-4.8 }{ 6 }=\frac{ 11.5 }{ 6 }=1.92 \approx.\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

u put it on 2 decimal places??

OpenStudy (anonymous):

much better now hehehe thanks you very much...

OpenStudy (anonymous):